Enrique cycled to his grandmother's house at 15 kph and returned home over the same route at 18 kph. If the whole trip took 7 hrs. and 20 min., how far does Enrique live from his grandmother's house?



Pls. ans. this..

Found 2 solutions by stanbon, checkley79:
Answer by stanbon(75887)

Enrique cycled to his grandmother's house at 15 kph and returned home over the same route at 18 kph. If the whole trip took 7 hrs. and 20 min., how far does Enrique live from his grandmother's house?
-----------------
To Grandmother's DATA:
rate = 15 kph ; distance = x km ; time = x/15 hrs
----------------
From Grandmother's DATA:
rate = 18 kph ; distance = x km ; time = x/18 hrs
-----------------------
Equation:
time + time = 7 1/3 hrs.
----
x/15 + x/18 = 22/3
----
Multiply thru by 90 to get:
6x + 5x = 660
11x = 660
x = 60 km (distance to Grandmother's)
========================================
cheers,
Stan H.

Answer by checkley79(3341)

D=RT
D=13X
D=18(7.33-X)
13X=18(7.33-X)
13X=131,94-18X
13X+18X=131.94
31X=131.94
X=`3`.94/31
X=4.256 HOURS FOR THE 13 MPH TRIP.
D=13*4.256
D=55.33 MILES ANS.
PROOF:
55.33=18(7.33-4.256)
55.33=18*3.074
55.33=55.33